integrate the following integral
∫ cos^2(x/2) dx

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int cos^{2}\frac{x}{2}dx
I = \frac{1}{2}\int (1+cosx)dx
I = \frac{1}{2}[\int dx+\int cosxdx]
I = \frac{1}{2}[x+\int cosxdx]
I = \frac{1}{2}[x+sinx]+c